# Gradient-Sign Masking for Task Vector Transport Across Pre-Trained Models

### v1

@@ Abstract
When a new release of a foundation model is published, practitioners typically need to repeat full fine-tuning, even if the same task has already been solved in the previous version. A promising alternative is to reuse the parameter changes (i.e., task vectors) that capture how a model adapts to a specific task. However, they often fail to transfer across different pre-trained models due to their misaligned parameter space. In this work, we show that the key to successful transfer lies in the sign structure of the gradients of the new model. Based on this insight, we propose GradFix, a novel method that approximates the ideal gradient sign structure and leverages it to transfer knowledge using only a handful of labeled samples. Notably, this requires no additional fine-tuning: the adaptation is achieved by computing a few gradients at the target model and masking the source task vector accordingly. This yields an update that is locally aligned with the target loss landscape, effectively rebasing the task vector onto the new pre-training. We provide a theoretical guarantee that our method ensures first-order descent. Empirically, we demonstrate significant performance gains on vision and language benchmarks, consistently outperforming naive task vector addition and few-shot fine-tuning.
@@ 1 Introduction
Over the past few years, the paradigm in deep learning has shifted from training models from scratch to fine-tuning large pre-trained models. Adapting these large models to downstream tasks is advantageous, as it leads to stronger performance at a fraction of the cost. Such a shift has been evident in natural language processing and computer vision, where pre-trained models such as BERT (Devlin et al., 2019), CLIP (Radford et al., 2021), and their successors (OpenAI et al., 2023; Liu et al., 2023) have become the standard starting point for developing new applications.
@@
Since companies and researchers often update the checkpoints by using more data or improved training pipelines, practitioners often face the need to repeat fine-tuning on the same downstream tasks. This creates redundancy: the work invested in adapting one release is not directly reusable on the next. To address this issue, several lines of research have investigated how to systematically relate or transfer knowledge across parameter spaces. The model rebasin literature (Ainsworth et al., 2023; Rinaldi et al., 2025) investigates how to align and merge independently trained models by exploiting permutation symmetries in their parameters. In parallel, task arithmetic (Ilharco et al., 2023; Ortiz-Jiménez et al., 2023; Yadav et al., 2023; Panariello et al., 2025b; Marczak et al., 2025) has shown that task vectors (i.e., the difference $\tau=\theta^{ft}-\theta^{0}$ between the base and fine-tuned parameters) can be added, subtracted, and merged across models to induce new capabilities. On a similar note, the literature on mode connectivity (Garipov et al., 2018; Frankle et al., 2020) demonstrates that different fine-tuned solutions can be linked by low-loss paths, highlighting that model parameters encode highly structured and transferable representations. Together, these advances suggest that parameters encode rich and transferable structure that can be systematically manipulated to obtain the desired behavior at reduced cost.
@@
In particular, Rinaldi et al. (2025) formalizes this setting and proposes a technique to transport task vectors across transformer-based architectures. However, there remains a substantial gap between the transported fine-tune and an actual fine-tuned model. This gap highlights a key challenge: while task vectors are informative about adaptation, their direct transfer across different pre-trains is not guaranteed to align with the loss geometry of the target model. In fact, naive transfer may introduce harmful directions in parameter space, i.e., components of the task vector that are misaligned with the descent directions of the target loss, thus increasing the loss and limiting its effectiveness. Addressing this problem is crucial both for reducing the cost of adapting rapidly evolving foundation models and for enabling their use in low-data regimes, where re-running full fine-tuning is infeasible.
@@
In this work, we introduce a framework for transporting task-specific knowledge across pre-trained models using gradient-sign masking. Our key insight is that, although a fine-tuning trajectory encodes valuable task information, its effectiveness on a new pre-trained model depends on the local loss geometry of the target. Inspired by findings from the optimization and distributed training literature (Bernstein et al., 2018; Alistarh et al., 2017), we exploit the observation that the sign of the gradient provides a robust surrogate for the descent direction. Leveraging this insight, we introduce a simple yet effective method to transport a task vector from a source model to a target pre-trained model: we mask the source task vector using the gradient signs of the target, keeping only the components aligned with the target’s local loss landscape. We further provide a formal guarantee that, to first order, this transported update reduces the target loss, ensuring a principled safeguard against harmful or misaligned transfer.
@@
Empirically, we show that this method enables highly effective transfer of fine-tuning knowledge from an outdated pre-trained model to a newer one, even in the low-data regime where gradients can only be estimated from a handful of samples, partially closing the gap between naive transfer and full fine-tuning on the target model. Our contributions are:
@@
  • Building on this insight, we propose GradFix, a simple yet theoretically grounded mechanism that filters the source task vector according to the local loss geometry of the target model. We formally prove that this guarantees, to first order, that the transported update reduces the target loss.
  • We demonstrate empirically that our method enables effective transport of fine-tuning knowledge across pre-trained models in both vision and text domains, even in the low-data regime where gradients must be estimated from only a handful of samples. This shows the practicality of our approach in scenarios where re-finetuning is infeasible.
2 Related Works
Model merging. A growing literature explores how to merge multiple fine-tuned checkpoints derived from the same pretrained model. Wortsman et al. (2022) introduced model soups, showing that simple weight averaging of fine-tuned checkpoints often improves generalization. Task arithmetic formalizes fine-tuning deltas as task vectors, which can be added or negated to edit model behavior (Ilharco et al., 2023). Building on this view, Yadav et al. (2023) proposed TIES-Merging, which resolves conflicts among task vectors by enforcing sign consistency before aggregation. For improved robustness and scalability, ATM interleaves tuning with merging (Zhou et al., 2024), and task-vector-based cluster vectors into compact representations (Zeng et al., 2024). Extending this line of research, curvature-aware composition improves merging and incremental adaptation (Porrello et al., 2025), while embedding recomposition enhances flexibility and transfer across tasks (Panariello et al., 2025a).
Model rebasin. A different family of methods focuses on explicit rebasin, aligning independently pretrained models into a shared parameterization so that task vectors can be transported across different pretrainings. Git Re-Basin introduced permutation matching to map two networks into a common basin (Ainsworth et al., 2023). For transformers, Imfeld et al. (2024) applies Optimal Transport to softly align components, while Rinaldi et al. (2025) proposes permutation- and spectral-based procedures that enable task vector transfer across distinct pretrained models. Extensions such as permutation least-squares (Nasery et al., 2025) and supernet formulations (Stoica et al., 2024) address more heterogeneous or large-scale settings.
Gradient information. A complementary line of work studies the utility of gradient signs and compressed gradient information. SignSGD (and its majority-vote variant) shows that one-bit sign information can suffice for convergence in distributed settings (Bernstein et al., 2018), while quantization and error-feedback analyses formalize guarantees for compressed updates (Alistarh et al., 2017; Karimireddy et al., 2019). More recent methods leverage gradient magnitudes or sign statistics for efficient adaptation: Gradient-Mask Tuning masks low-importance parameters during LLM fine-tuning (Li et al., 2025), and sign-based federated variants weight client updates to address heterogeneity (Park et al., 2024).
@@ 3 Preliminaries
Let $\theta_{A}$ and $\theta_{B}$ denote the parameters of the same architecture, pre-trained on different datasets (or with different hyperparameters). The fine-tune of $\theta_{A}$ on a downstream dataset $\mathcal{D}$ is denoted as $\theta_{A}^{ft}$.
Model rebasin. The goal of rebasin (Ainsworth et al., 2023) is to align two independently trained models by mapping the parameters of one into the loss basin of the other, so that they become functionally compatible. This setting concerns only the pre-trained weights, without involving fine-tuning updates.
Task Arithmetic. A complementary perspective to model rebasin is offered by task arithmetic (Ilharco et al., 2023; Yadav et al., 2023), which studies linear operations on task-specific parameter updates. Given a pre-trained model $\theta^{0}$ and a fine-tuned counterpart $\theta^{\star}$, the difference vector in parameter space $\tau=\theta^{\star}-\theta^{0}$ is called a task vector. Task vectors describe how a base model adapts to the task and can be added, subtracted, or merged to induce new behaviors. This setting usually assumes that all models share the same initialization $\theta^{0}$, which ensures comparability across tasks.
Our Setting. In contrast, our goal is to apply task-vector style transfer when the same base initialization assumption does not hold: we want to transfer $\tau_{A}=\theta_{A}^{ft}-\theta_{A}$ from a source pre-train $\theta_{A}$ onto a different pre-train $\theta_{B}$. This connects to rebasin in that the bases differ, but unlike traditional rebasin approaches, we do not seek to explicitly align parameterizations. Instead, we ask:
@@ 4 Method
GradFix is a framework for transferring task vectors across different pre-trains by filtering them with gradient information from the target model. As a conceptual starting point, we consider an oracle ideal setting where the target task vector, obtained from full fine-tuning, defines the ideal transferable directions (Sec. 4.1). Then the oracle is approximated with a single gradient step on the target model, using the signs of the gradients to capture an approximate direction of the full fine-tuning trajectory. This yields a gradient-sign mask that selectively filters the source task vector into a compatible update (Sec. 4.2). Finally, we extend the approach to the limited-data regime, where gradients are estimated from only a handful of labeled samples (Sec. 4.3).
@@ 4.1 GradFix (Gradient-Sign Masking)
We begin by considering an ideal scenario where the true fine-tuned task vector $\tau_{B}:=\theta_{B}^{ft}-\theta_{B}$ of the model $B$ and the whole target dataset $\mathcal{D}$ are available. Such a vector represents the optimal parameter change to adapt $B$ on the target dataset $\mathcal{D}$. In this ideal setting, it is possible to construct a mask that perfectly retains only the components of a candidate update (e.g., $\tau_{A}$) that are aligned with $\tau_{B}$, ensuring that every retained coordinate contributes to decreasing the loss. In other words, $\tau_{B}$ (or its sign structure) defines the “gold standard” for locally beneficial directions. Formally, we define as ${\bm{m}}^{\star}\in\{0,1\}^{d}$ the mask induced by $\tau_{B}$, where $d$ is the total number of model parameters and $i\in\{1,\dots,d\}$ indexes each coordinate:
$m^{\star}_{i}=\mathbbm{1}\{\mathrm{sign}(\tau_{A,i})=\mathrm{sign}(\tau_{B,i})\}.$ (1)
@@
where $\odot$ denotes the element-wise multiplication. This vector $\delta^{\star}$ represents a reliable transfer of $\tau_{A}$ onto $\theta_{B}$, since it filters out all components of $\tau_{A}$ that are misaligned with the true adaptation directions of $B$. In practice, however, $\tau_{B}$ (and thus $\delta^{\star}$) is unavailable because it requires access to the fine-tuned target model $\theta_{B}^{ft}$, which defeats the purpose of transporting the solution from $A$ to $B$. To approximate this ideal mask, we consider the gradient of the zero-shot target model as a surrogate for $\tau_{B}$, indicating locally beneficial directions:
@@
where $\ell$ is the training objective (e.g., cross-entropy) and $(x,y)$ is a labeled example from $\mathcal{D}$. Based on this gradient, we define the gradient-sign mask $\bm{m}$}, which retains only the components of $\tau_{A}$ whose sign matches that of the corresponding gradient coordinate:
$m_{i}:=\mathbbm{1}\{\mathrm{sign}(\tau_{A,i})=\mathrm{sign}(g_{i})\}.$ (4)
Intuitively, $\bm{g}$} acts as a signal for local alignment with the loss geometry of $B$. Notably, in the idealized setting where $B$ is fine-tuned using full-batch gradient descent for a single epoch, the resulting task vector $\tau_{B}$ is exactly proportional to $\bm{g}$}, so the gradient-sign mask coincides with the oracle. This observation provides a clear justification for using the gradient-sign mask as an approximation of the ideal update, even when only a few labeled examples are available. The gradient-sign mask selectively retains only the components of $\tau_{A}$ whose sign matches the gradient of $\mathcal{L}(\theta_{B})$, effectively pruning coordinates that would increase the loss for the target model $B$. In this way, the gradient-sign mask provides a practical surrogate for the trajectory-informed directions encoded in the unavailable $\tau_{B}$, capturing the locally beneficial update directions without requiring access to the fully fine-tuned target model.
@@ 4.2 Transporting the Update
$\theta_{B}^{\text{trans}}=\theta_{B}-\delta^{A},\quad\delta^{A}:=\alpha\,({\bm{m}}\odot\tau_{A}),$ (5)
It is important to note that $\tau_{A}$ points in a descent direction for model $A$, whereas the gradient $\bm{g}$} of the target model points in the ascent direction of its loss. By subtracting the masked vector $\delta^{A}$, we ensure that each retained component moves opposite to the ascent direction of $B$, producing a descent-aligned update, whereas $\delta^{\star}$ already points in the descent direction since it is computed directly from $\tau_{B}$.
Descent guarantee. To understand why such a gradient masking provides effective transfer, we analyze its effect on the loss of the target model $B$. Consider the transported update from Eq. 5, by expanding the target loss $\mathcal{L}$ around $\theta_{B}$ via a first-order Taylor approximation we obtain:
$\mathcal{L}(\theta_{B}-\alpha\delta^{A})\approx\mathcal{L}(\theta_{B})-\alpha{\bm{g}}^{\top}\delta^{A},\quad\text{where}\ \ {\bm{g}}=\nabla_{\theta}\mathcal{L}(\theta_{B}).$ (6)
@@
$g_{i}\cdot(m_{i}\tau_{A,i})=m_{i}|g_{i}|\,|\tau_{A,i}|\;\geq\;0,$ (7)
so that the overall inner product satisfies the following:
$\alpha{\bm{g}}^{\top}\delta^{A}=\alpha\sum_{i}m_{i}|g_{i}|\,|\tau_{A,i}|\;\geq\;0.$ (8)
Thus, this implies that, for sufficiently small $\alpha$, the update $\delta^{A}$ is guaranteed to be a descent direction for $\mathcal{L}$. Practically, the mask removes all sign-mismatched components of $\tau_{A}$, so that every retained entry contributes positively to reducing the loss. Without masking, $\tau_{A}$ could contain harmful directions that increase the loss for $B$; with masking, the transported update is locally aligned with the descent geometry of the target model.

### v3

@@ Abstract
When a new release of a foundation model is published, practitioners typically need to repeat fine-tuning, even if the same task was already tackled in the previous version. A promising alternative is to reuse the parameter changes (i.e., task vectors) that capture how a model adapts to a specific task. However, these vectors often fail to transfer across different pre-trained models because their parameter spaces are misaligned. In this work, we show that successful transfer depends strongly on the gradient-sign structure of the new model. Based on this insight, we propose GradFix, which approximates the ideal sign structure and leverages it to transfer knowledge using only a handful of labeled samples. Notably, this requires no additional fine-tuning: we only compute a few target-model gradients without parameter updates and mask the source task vector accordingly. This yields an update that is locally aligned with the target loss landscape, effectively rebasing the task vector onto the new pre-training. We provide a theoretical guarantee that our method ensures first-order descent. Empirically, we demonstrate significant performance gains on vision and language benchmarks, consistently outperforming naive task vector addition and few-shot fine-tuning. We further show that transporting task vectors improves multi-task and multi-source model merging. Code is available at https://github.com/fillo-rinaldi/GradFix.
@@ 1 Introduction
Over the past few years, the paradigm in deep learning has shifted from training models from scratch to fine-tuning large pre-trained models. Adapting these large models to downstream tasks is advantageous, as it leads to stronger performance at a fraction of the cost. Such a shift has been evident in natural language processing and computer vision, where pre-trained models such as BERT (devlin2019bert), CLIP (radford2021learning), and their successors (openai2023gpt04; liu2023llava) have become the standard starting point for developing new applications.
@@
Since companies and researchers often update checkpoints using more data or improved training pipelines, practitioners frequently need to repeat fine-tuning on the same downstream tasks. This creates redundancy: the work invested in adapting one release is not directly reusable on the next. To address this issue, several lines of research have investigated how to systematically relate or transfer knowledge across parameter spaces. The model rebasin literature (ainsworth2023git; rinaldi2025update) studies how to align and merge independently trained models by exploiting permutation symmetries in their parameters. In parallel, Task Arithmetic (ilharco2023task; ortizjimenez2023tangent; yadav2023tiesmerging; panariello2025accurate; porrello2026dataless) has shown that task vectors (i.e., the difference $\tau=\theta^{ft}-\theta^{0}$ between base and fine-tuned parameters) can be added, subtracted, and merged across models to induce new capabilities. On a similar note, the literature on mode connectivity (garipov2018loss; frankle2020linear) demonstrates that different fine-tuned solutions can be linked by low-loss paths, highlighting that model parameters encode highly structured and transferable representations. Together, these advances suggest that parameters encode rich and transferable structure that can be systematically manipulated to obtain the desired behavior at reduced cost.
@@
In particular, rinaldi2025update formalizes this setting and proposes a technique to transport task vectors across transformer-based architectures. Yet a large gap remains between the transported fine-tune and an actual fine-tuned model, which highlights a key challenge: while task vectors are informative about adaptation, their direct transfer across different pre-trained models is not guaranteed to align with the loss geometry of the target model. In fact, naive transfer may introduce harmful directions in parameter space, i.e., components of the task vector that are misaligned with the descent directions of the target loss, thus increasing the loss and limiting its effectiveness. Addressing this problem is crucial both for reducing the cost of adapting rapidly evolving foundation models and for enabling their use in low-data regimes, where re-running full fine-tuning is infeasible.
@@
In this work, we introduce a framework for transporting task-specific knowledge across pre-trained models using gradient-sign masking. Our key insight is that, although a fine-tuning trajectory encodes valuable task information, its effectiveness on a new pre-trained model depends on the local loss geometry of the target. Inspired by findings from the optimization and distributed training literature (bernstein2018signsgd; alistarh2017qsgd), we exploit the observation that the sign of the gradient provides a robust surrogate for the descent direction. Leveraging this insight, we introduce a simple yet effective method to transport a task vector from a source model to a target pre-trained model: we mask the source task vector using the gradient signs of the target, keeping only the components aligned with the target’s local loss landscape. We further provide a formal guarantee that, to first order, this transported update reduces the target loss, ensuring a principled safeguard against harmful or misaligned transfer.
@@
Empirically, we show that this method enables highly effective transfer of fine-tuning knowledge from an outdated pre-trained model to a newer one, even in the low-data regime where gradients can only be estimated from a handful of samples, partially closing the gap between naive transfer and full fine-tuning on the target model. We also evaluate GradFix in model-merging pipelines, showing gains in both multi-task and multi-source transport settings. Our contributions are:
@@
  • Building on this insight, we propose GradFix, a simple mechanism that filters the source task vector using the target model’s local loss geometry. We formally prove that, to first order, the transported update reduces the target loss.
  • We empirically show that our method enables effective transport of fine-tuning knowledge across pre-trained models in both vision and text domains, even in the low-data regime where gradients must be estimated from only a handful of samples. We further validate that GradFix improves model-merging performance in both multi-task and multi-source settings, showing that the transported updates remain useful beyond single-task transfer.
2 Related Work
Model merging. Prior work studies how to merge fine-tuned checkpoints from the same pre-trained model. Model soups show that simple weight averaging can improve generalization (wortsman2022model). Task Arithmetic interprets fine-tuning deltas as task vectors that can be combined through linear operations (ilharco2023task), and TIES-Merging addresses conflicts by enforcing sign consistency (yadav2023tiesmerging). Recent extensions include curvature-aware composition for improved adaptation (porrello2025a) and modular embedding recomposition for more flexible transfer across diverse tasks (panariello2025modular).
Model rebasin. A different family of methods focuses on explicit rebasing, aligning independently pre-trained models into a shared parameterization so that task vectors can be transported across different pre-trainings. Git Re-Basin introduced permutation matching to map two networks into a common basin (ainsworth2023git). For transformers, imfeld2024transformer applies Optimal Transport to softly align components, while rinaldi2025update proposes permutation- and spectral-based procedures that enable task vector transfer across distinct pre-trained models. Extensions such as permutation least-squares (nasery2025pleas) and supernet formulations (stoica2024zipit) address more heterogeneous or large-scale settings.
Gradient information. A complementary line of work studies the utility of gradient signs and compressed gradient information. SignSGD (and its majority-vote variant) shows that one-bit sign information can suffice for convergence in distributed settings (bernstein2018signsgd), while quantization and error-feedback analyses formalize guarantees for compressed updates (alistarh2017qsgd; karimireddy2019error). More recent methods leverage gradient magnitudes or sign statistics for efficient adaptation: Gradient-Mask Tuning masks low-importance parameters during LLM fine-tuning (li2025enhancing), and sign-based federated variants weight client updates to address heterogeneity (park2024signsgd).
@@ 3 Preliminaries
Let $\theta_{A}$ and $\theta_{B}$ denote the parameters of the same architecture, pre-trained on different datasets (or with different hyperparameters). The fine-tuned $\theta_{A}$ on a downstream task is denoted by $\theta_{A}^{ft}$.
Model rebasin. The goal of rebasin (ainsworth2023git) is to align two independently trained models by mapping the parameters of one into the loss basin of the other, so that they become functionally compatible. This setting concerns only the pre-trained weights, without involving fine-tuning updates.
Task Arithmetic. A complementary perspective to model rebasin is offered by Task Arithmetic (ilharco2023task; yadav2023tiesmerging), which studies linear operations on task-specific parameter updates. Given a pre-trained model $\theta^{0}$ and a fine-tuned counterpart $\theta^{\star}$, the difference vector in parameter space $\tau:=\theta^{\star}-\theta^{0}$ is called a task vector. Task vectors describe how a base model adapts to the task and can be added, subtracted, or merged to induce new behaviors. This setting usually assumes that all models share the same initialization $\theta^{0}$, which ensures comparability across tasks.
Our setting. In contrast, our goal is to apply task-vector transfer when the target and source models do not share the same initialization: we want to transfer $\tau_{A}=\theta_{A}^{ft}-\theta_{A}$ from a source pre-trained model $\theta_{A}$ to a different pre-trained model $\theta_{B}$. This connects to rebasin in that the bases differ, but unlike traditional rebasin approaches, we do not seek to explicitly find a permutation to align parameters. Instead, we ask:
@@ 4 Method
Here we introduce GradFix , a framework for transferring task vectors across different pre-trained models by filtering them with gradient information from the target model. As a conceptual starting point, we first consider an oracle setting where the target task vector obtained from full fine-tuning defines the ideal transferable directions (Sec. 4.1). We then approximate this oracle with a single gradient step on the target model, using anti-gradient signs to estimate the direction of the full fine-tuning trajectory. This yields a gradient-sign mask that selectively filters the source task vector into a compatible update (Sec. 4.2). Finally, we extend the approach to the limited-data regime, where gradients are estimated from only a handful of labeled samples (Sec. 4.3).
@@ 4.1 GradFix (Gradient-Sign Masking)
We begin by considering an ideal scenario where the true fine-tuned task vector $\tau_{B}=\theta_{B}^{ft}-\theta_{B}$ of model $B$ and the full target dataset $\mathcal{D}$ are available. This vector represents the optimal parameter change to adapt $B$ to $\mathcal{D}$. In this setting, we can construct a mask that retains only the components of a candidate update (e.g., $\tau_{A}$) that are aligned with $\tau_{B}$, ensuring that every retained coordinate contributes to decreasing the loss. In other words, $\tau_{B}$ (or its sign structure) defines the gold standard for locally beneficial directions. Formally, we define the oracle mask ${\bm{m}}^{\star}\in{\{0,1\}}^{d}$, where $d$ is the total number of model parameters and $i\in\{1,\dots,d\}$ indexes a parameter coordinate:
$m^{\star}_{i}=\mathbbm{1}\{\operatorname{sign}(\tau_{A,i})=\operatorname{sign}(\tau_{B,i})\}.$ (1)
@@
where $\odot$ denotes element-wise multiplication. This vector $\delta^{\star}$ represents a reliable transfer of $\tau_{A}$ onto $\theta_{B}$, since it filters out all components of $\tau_{A}$ that are misaligned with the true adaptation directions of $B$. In practice, however, $\tau_{B}$ (and thus $\delta^{\star}$) is unavailable because it requires access to the fine-tuned target model $\theta_{B}^{ft}$, which defeats the purpose of transporting the solution from $A$ to $B$. To approximate this ideal mask, we use the gradient of the zero-shot target model as a surrogate for $\tau_{B}$, since it captures locally beneficial directions:
@@
where $\ell$ is the training objective (e.g., cross-entropy) and $(x,y)$ is a labeled example from $\mathcal{D}$. Based on this gradient, we define the gradient-sign mask $\bm{m}$}, which retains only the components of $\tau_{A}$ whose sign matches that of the corresponding anti-gradient coordinate:
$m_{i}:=\mathbbm{1}\{\operatorname{sign}(\tau_{A,i})=\operatorname{sign}(-g_{i})\}.$ (4)
Intuitively, $\bm{g}$} acts as a signal for local alignment with the loss geometry of $B$. Notably, in the idealized setting where $B$ is fine-tuned using full-batch gradient descent for a single epoch, the resulting task vector $\tau_{B}$ is proportional to $\bm{g}$}, so the gradient-sign mask coincides with the oracle. This observation justifies using the gradient-sign mask as an approximation of the ideal update, even when only a few labeled examples are available. The mask retains only components of $\tau_{A}$ whose sign matches the anti-gradient of $\mathcal{L}(\theta_{B})$, pruning coordinates that would increase the loss for target model $B$. In this way, the gradient-sign mask serves as a practical surrogate for the trajectory-informed directions encoded in the unavailable $\tau_{B}$.
@@ 4.2 Transporting the Update
$\theta_{B}^{\text{trans}}=\theta_{B}+\delta^{A},\quad\delta^{A}:=\alpha\,({\bm{m}}\odot\tau_{A}).$ (5)
We explicitly use the convention ${\bm{g}}=\nabla_{\theta}\mathcal{L}(\theta_{B})$ (ascent direction), so useful transport directions should align with $\bm{g}$}. It is important to note that $\tau_{A}$ points in a descent direction for model $A$, whereas the gradient $\bm{g}$} of the target model points in the ascent direction of its loss. By selecting only coordinates aligned with $\bm{g}$} and then adding $\delta^{A}$, each retained component moves along a descent-aligned direction for $B$. In contrast, $\delta^{\star}$ is oracle-aligned with the target task direction because its mask is constructed from $\tau_{B}$.
This construction induces a coordinate-wise filtering effect: we do not alter the direction of retained entries, but only suppress coordinates that are sign-incompatible with the target anti-gradient. As a result, the transported update preserves task information from $\tau_{A}$ while reducing the risk of injecting locally harmful directions into $\theta_{B}$. We now formalize this intuition with a first-order loss analysis.
Descent guarantee. To understand why this gradient masking provides effective transfer, we analyze its effect on the loss of the target model $B$. Consider the transported update from Eq. 5. By expanding the target loss $\mathcal{L}$ around $\theta_{B}$ via a first-order Taylor approximation, we obtain:
$\mathcal{L}(\theta_{B}+\delta^{A})\approx\mathcal{L}(\theta_{B})+{\bm{g}}^{\top}\delta^{A},\quad\text{where}\ \ {\bm{g}}=\nabla_{\theta}\mathcal{L}(\theta_{B}).$ (6)
@@
$g_{i}\cdot(m_{i}\tau_{A,i})=\begin{cases}-|g_{i}|\,|\tau_{A,i}|,&\text{if }\operatorname{sign}(\tau_{A,i})=\operatorname{sign}(-g_{i}),\\ 0,&\text{otherwise},\end{cases}$ (7)
which is always nonpositive. Coordinates with $g_{i}=0$ or $\tau_{A,i}=0$ are naturally covered by this expression and contribute zero. Therefore, the overall inner product satisfies the following:
${\bm{g}}^{\top}\delta^{A}=-\alpha\sum_{i}m_{i}|g_{i}|\,|\tau_{A,i}|\;\leq\;0.$ (8)
Thus, for sufficiently small $\alpha$, the update $\delta^{A}$ is guaranteed to be a descent direction for $\mathcal{L}$. Practically, the mask removes all sign-mismatched components of $\tau_{A}$, so that every retained entry contributes to reducing the loss. Without masking, $\tau_{A}$ could contain harmful directions that increase the loss for $B$; with masking, the transported update is locally aligned with the descent geometry of the target model.